# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 5.2.2 Second transformation, model problem and its solution: the case $\mathrm{Re}\,(s)<0$
Solution of RHP 5.12.
With the same definition of $R$ in (5.60), we also set the same definition of $h$ as in (5.84) where however now the domain consists of ${\mathbb{C}}$ minus the branch cuts $\Gamma_{b}^{1}\cup\Gamma_{b}^{a_{+}}\cup\Gamma_{a_{+}}^{a_{-}}$, and the determination chosen so that $h(z)\simeq\frac{1}{z}$ as $z\to\infty$. A careful analysis of the phases of $h$ shows that
$h(z_{+})$ $=-ih(z_{-})$ $z\in\Gamma_{b}^{1}$
$h(z_{+})$ $=-ih(z_{-})$ $z\in\Gamma_{a_{+}}^{a_{-}}$
$h(z_{+})$ $=-h(z_{-})$ $z\in\Gamma_{b}^{a_{+}}.$ (5.123)
@@
Lemma 5.2 (Properties of the Abel map)
@@
$\mathfrak{u}(z_{+})=-\mathfrak{u}(z_{-})+\left\{\begin{array}[]{lc}0&z\in\Gamma_{a_{+}}^{a_{-}}\\ -\tau&z\in\Gamma_{b}^{1}\end{array}\right.$ (5.126)
$\mathfrak{u}(z_{+})=\mathfrak{u}(z_{-})-1,\qquad\qquad z\in\Gamma_{b}^{a_{+}}.$ (5.127)
Consider now the following two row-vectors;
$\boldsymbol{\phi}(z)$ $=[\phi_{1}(z),\phi_{2}(z)],$
$\phi_{1}(z)$ $=\frac{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{i\pi K\mathfrak{u}(z)}$
$\phi_{2}(z)$ $=\frac{i\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-i\pi K\mathfrak{u}(z)}$ (5.128)
@@
$\boldsymbol{\psi}(z)$ $=[\psi_{1}(z),\psi_{2}(z)]$
$\psi_{1}(z)$ $=\frac{\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{i\pi K\mathfrak{u}(z)}$
$\psi_{2}(z)$ $=\frac{i\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-i\pi K\mathfrak{u}(z)}.$ (5.129)
These are essentially the same formulæ as (5.95) (5.96) except for a minor modification of the normalization. Accordingly we define $\widehat{M}(z;G,K,\infty)$ exactly as in (5.98). The boundary relations for this matrix $\widehat{M}$ are now:
$\widehat{M}(z_{+})$ $=\widehat{M}(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ $z\in\Gamma_{a_{+}}^{a_{-}}$
$\widehat{M}(z_{+})$ $=\widehat{M}(z_{-})\begin{bmatrix}-{\rm e}^{-i\pi K}&0\\ 0&-{\rm e}^{i\pi K}\end{bmatrix}$ $z\in\Gamma_{b}^{a_{+}}$
$\widehat{M}(z_{+})$ $=\widehat{M}(z_{-})\begin{bmatrix}0&{\rm e}^{-2i\pi\left(G-\frac{K\tau}{2}\right)}\\ -{\rm e}^{2i\pi\left(G-\frac{K\tau}{2}\right)}&0\end{bmatrix}$ $z\in\Gamma_{b}^{1}.$ (5.130)
By matching these to the boundary relations (5.122) we deduce the following:
Proposition 5.5
The solution of the RHP (5.12) is given by $\widehat{M}(z;G,K)$ in (5.98) with the values of the constants $G,K$ given by
$G=\frac{1}{2i\pi}\left(n\Omega_{1}+\nu+n\Omega_{2}\tau\right)+\frac{\tau}{2},\qquad K=\frac{n\Omega_{2}}{i\pi}+1.$ (5.131)
The proof is entirely similar to Prop. 5.3. The same observation holds that the solvability of the RHP 5.12 and hence of the RHP 5.12 depends entirely on the non-vanishing of the expression
$\vartheta\left(\frac{n\Omega_{1}+\nu-n\Omega_{2}\tau}{2i\pi}+\frac{1}{2}\right)$ (5.132)
@@
$\widetilde{H}_{1}:=\frac{n\Omega_{1}}{2i\pi}+\mathrm{Re}\,\left(\frac{\nu}{2i\pi}\right)-\frac{\mathrm{Re}\,\tau}{\mathrm{Im}\,\tau}\mathrm{Im}\,\left(\frac{\nu}{2i\pi}\right)$ $\in{\mathbb{Z}}$
$\widetilde{H}_{2}:=\frac{n\Omega_{2}}{2i\pi}+\frac{\mathrm{Im}\,\left(\frac{\nu}{2i\pi}\right)}{\mathrm{Im}\,\tau}+\frac{1}{2}$$\in{\mathbb{Z}}.$ (5.133)
The mesh of level sets $\widetilde{H}_{1}\in{\mathbb{Z}}\ni\widetilde{H}_{2}$ are the gridlines shown in Fig. 3 in the portion $\mathrm{Re}\,(s)<0$ of each pane.
@@ 5.2.3 Local parametrices
The construction of the local parametrices is similar to that of Section 5.1.1. The only parametrix that is less commonly encountered (but see [10]) is the one that needs to be defined at the point $b$ in the case $\mathrm{Re}\,(s)>0$, when three main arcs meet at the same point (see Fig. 8, left pane). Let $\mathbb{D}_{x}$ denote a small disk centered at $x\in\{a_{+},a_{-},b,1\}$ and not containing any of the other branchpoints $a_{+},a_{-},b,1$. Inside each of these disks one has to define an appropriate local solution of the RHP 5.7 or 5.10 (according to the sign of $\mathrm{Re}\,(s)>0$ or $\mathrm{Re}\,(s)<0$, respectively). Such a local solution (a.k.a. “local parametrix”) is constructed in terms of Airy functions or Bessel functions. More specifically:
  1. 1.
@@
  2. 2.
    Inside $\mathbb{D}_{b}$ the local parametrix is constructed also in terms of Airy functions, arranged however in a different formula along the lines explained in App. A.1.
  3. 3.
    Inside $\mathbb{D}_{1}$ the local parametrix is constructed in terms of Bessel functions; this can be found in ([58], Section 3, pag. 155–156), or ([8], Section 4.2), and we will not enter into details since it is mainly irrelevant to the formulas that are of interest in this paper.
5.2.4 Summary; the approximating mesh of location of zeros
@@
$J_{\mathscr{E}}(z)=M_{{}_{Q}}(z)\left(\mathbf{1}+\mathcal{O}(n^{-1})\right)M_{{}_{Q}}^{-1}(z).$ (5.134)
While in the region outside $EoT$ the model solution $M_{{}_{Q}}$ does not depend on $n$ (see (5.1)) and hence $J_{\mathscr{E}}=\mathbf{1}+\mathcal{O}(n^{-1})$ since the matrix $M_{{}_{Q}}$ is $n$-independent . On the other hand when $s\in EoT$ the solution $M$ is instead given by Prop. 5.3 and consequently $M_{{}_{Q}}$ (5.80)) depends on $n$.
The validity of the estimate $J_{\mathscr{E}}=\mathbf{1}+\mathcal{O}(n^{-1})$ depends on the matrices $M$, $M^{-1}$ remaining bounded, uniformly in $n$, with respect to $z$ on the contour. Given that $\det(M)\equiv 1$, the source of the potential unboundedness are the terms in the denominators in (5.98). Of these the only term that may potentially vanish is the term $\vartheta\left(G-\frac{\tau+1}{2}\right)$, which becomes (5.103) or (5.132) (depending on the sign of $\mathrm{Re}\,(s)$). Therefore the correct error estimate is
$J_{\mathscr{E}}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{n\vartheta\left(G-\frac{\tau+1}{2}\right)^{2}}\right),$ (5.135)
with $G$ given by either (5.101) or (5.131) (depending on the sign of $\mathrm{Re}\,(s)$).
Thus the sufficient condition of asymptotic solvability of the RHP 5.1 is that
$\vartheta\left(\frac{n\Omega_{1}(s)-n\Omega_{2}(s)\tau(s)+\nu(s)}{2i\pi}+\frac{1}{2}\mathcal{X}_{\{\mathrm{Re}\,s>0\}}(s)\right)=o\left(\frac{1}{\sqrt{n}}\right),$ (5.136)
with $\mathcal{X}_{\{\mathrm{Re}\,s>0\}}(s)$ denoting the indicator function of the right half–plane. This condition fails inside the union of small disks of radii $\mathcal{O}(n^{-\frac{3}{2}})$ in the $s$–plane centered at the points where $\vartheta=0$, namely, the intersection of the mesh of lines indicated by (5.2.1) (for $\mathrm{Re}\,(s)>0$) or (5.2.2) (for $\mathrm{Re}\,(s)<0$).
The actual numerics shows a seemingly much faster convergence of the mesh to the actual position of the zeros (see Fig. 3), which appears to be one of those serendipitous situations where the approximation works “better than expected”.
Figure 10: The RHPs for the error terms in the two cases $\mathrm{Re}\,(s)>0$ (left) and $\mathrm{Re}\,(s)<0$. Compare with Fig. 8.
@@ 6 Approximation of the Hamiltonian
The Hamiltonian is obtained from the Taylor expansion near $z=0$ of the solution of RHP 5.1 outside of $EoT$, and RHP 5.5 inside $EoT$, according to Proposition 2.2.
In either case the chain of transformations ($Y\to W\to T_{0}\sim M_{{}_{Q}}$) and the subsequent approximation imply that in a neighbourhood of $z=0$ we have
$Y_{n}(z)={\rm e}^{-n\frac{\ell}{2}\sigma_{3}}W(z){\rm e}^{n\left(g(z)+\frac{\ell}{2}\right)\sigma_{3}}={\rm e}^{-n\frac{\ell}{2}\sigma_{3}}\mathscr{E}(z)M_{{}_{Q}}(z){\rm e}^{n\left(g(z)+\frac{\ell}{2}\right)\sigma_{3}},$ (6.1)
where $\mathscr{E}(z)$, as discussed in Sec. 5.2.4, is a matrix that is $\mathbf{1}+\mathcal{O}\left(\frac{1}{n\vartheta\left(G-\frac{\tau+1}{2}\right)^{2}}\right)$ near the origin (as long as $s$ has finite distance from the boundary of $EoT$ and the imaginary axis within it).
@@
$Y_{n}^{-1}Y_{n}^{\prime}\bigg{|}_{z=0}=\left({\rm e}^{-ng\sigma_{3}}\left(M_{{}_{Q}}^{-1}M_{{}_{Q}}^{\prime}+M^{-1}_{{}_{Q}}\mathscr{E}^{-1}\mathscr{E}^{\prime}M_{{}_{Q}}\right){\rm e}^{ng\sigma_{3}}+ng^{\prime}(z)\sigma_{3}\right)\bigg{|}_{z=0}.$ (6.2)
Given the expression of $M$ in either Prop. 5.1 (for $s\not\in EoT$), Prop. 5.3 for $s\in EoT\cap\{\mathrm{Re}\,(s)>0\}$, or Prop. 5.5 $s\in EoT\cap\{\mathrm{Re}\,(s)<0\}$ we have
$Y_{n}^{-1}Y_{n}^{\prime}\bigg{|}_{z=0}=\left({\rm e}^{(-ng-S)\sigma_{3}}\left(M^{-1}M^{\prime}+M^{-1}\mathscr{E}^{-1}\mathscr{E}^{\prime}M\right){\rm e}^{(S+ng)\sigma_{3}}+\left(S^{\prime}+ng^{\prime}(z)\right)\sigma_{3}\right)\bigg{|}_{z=0},$ (6.3)
@@
We are interested in the $(1,1)$ entry and hence the conjugation in the first term by the diagonal matrix ${\rm e}^{(S-ng)\sigma_{3}}$ is immaterial. Furthermore the term containing $\mathscr{E}$ yields a sub-leading contribution which we ignore for the purpose of this computation. In principle we should fork the computation according to the three regions (i) outside $EoT$; (ii) inside $EoT$ with $\mathrm{Re}\,(s)>0$; (iii) inside $EoT$ with $\mathrm{Re}\,(s)<0$. However the computation inside $EoT$ would be rather formal because we have not specified the correct orders of approximations when $s$ is in a neighbourhood of one of the zeros. Additionally the explicit approximation is of no particular interest to us and thus for simplicity we decided to forego it entirely in this paper.
Hamiltonian outside $EoT$.
Using the explicit expression of $M(z)$ in Prop. 5.1 we have
$M^{-1}M^{\prime}=\frac{i\sigma_{2}}{s}$ (6.4)
and hence this term does not contribute being off–diagonal. Using (5.22) and the fact that $R(0)=-\frac{s}{2},\ \ R(1)=-\sqrt{1+\frac{s^{2}}{4}}$ with the root’s determination such that $R(1;s)\simeq-\frac{s}{2}$ as $|s|\to\infty$, we find
$S^{\prime}(0)=\frac{{K}}{s}+\frac{2{\rho}}{s^{2}+s\sqrt{s^{2}+4}}.$ (6.5)
Using then the explicit expression of $g(z)$ derived from (4.17)
$g^{\prime}(z)=\frac{1}{2}\varphi^{\prime}(z;s)+\frac{s}{2z^{2}}=\frac{\sqrt{z^{2}+\frac{s^{2}}{4}}}{z^{2}}+\frac{s}{2z^{2}}.$ (6.6)
Since $g^{\prime}(z)$ is regular at $z=0$ we need to take the root’s determination near $z=0$ that tends to $-\frac{s}{2}$ and hence
$g^{\prime}(z)=-s\frac{\sqrt{1+\frac{4z^{2}}{s^{2}}}}{2z^{2}}+\frac{s}{2z^{2}}=-\frac{1}{s}+\mathcal{O}(z)\ \ \ \Rightarrow\ \ \ g^{\prime}(0;s)=-\frac{1}{s}.$ (6.7)
We thus conclude that
Proposition 6.1
For $s$ in closed subsets outside $EoT$ we have the uniform approximation
$H_{V}(s)=\frac{n-{K}}{s}-\frac{2{\rho}}{s^{2}+s\sqrt{s^{2}+4}}+\frac{{\rho}}{2}+\mathcal{O}(n^{-1}).$ (6.8)

### v3

@@ 5.2.2 Second transformation, model problem and its solution: The case $\boldsymbol{\operatorname{Re}(s)<0}$
Solution of Riemann–Hilbert Problem 5.23. With the same definition of $R$ in (5.19), we also set the same definition of $h$ as in (5.30) where however now the domain consists of $\mathbb{C}$} minus the branch cuts $\Gamma_{b}^{1}\cup\Gamma_{b}^{a_{+}}\cup\Gamma_{a_{+}}^{a_{-}}$}, and the determination chosen so that $h(z)\simeq\frac{1}{z}$ as $z\to\infty$. A careful analysis of the phases of $h$ shows that
$h(z_{+})=-{\rm i}h(z_{-}),\qquad z\in\Gamma_{b}^{1},\qquad h(z_{+})=-{\rm i}h(z_{-}),\qquad z\in\Gamma_{a_{+}}^{a_{-}},$
$h(z_{+})=-h(z_{-}),\qquad z\in\Gamma_{b}^{a_{+}}.$
@@
Lemma 5.24 (properties of the Abel map).
@@
$\mathfrak{u}(z_{+})=-\mathfrak{u}(z_{-})+\begin{cases}0,&z\in\Gamma_{a_{+}}^{a_{-}},\\ -\tau,&z\in\Gamma_{b}^{1},\end{cases}\qquad\mathfrak{u}(z_{+})=\mathfrak{u}(z_{-})-1,\qquad z\in\Gamma_{b}^{a_{+}}.$
Consider now the following two row-vectors:
$\boldsymbol{\phi}(z)=[\phi_{1}(z),\phi_{2}(z)],\qquad\phi_{1}(z)=\frac{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{{\rm i}\pi K\mathfrak{u}(z)},$
$\phi_{2}(z)=\frac{{\rm i}\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)-\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-{\rm i}\pi K\mathfrak{u}(z)}$
@@
$\boldsymbol{\psi}(z)=[\psi_{1}(z),\psi_{2}(z)],\qquad\psi_{1}(z)=\frac{\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{{\rm i}\pi K\mathfrak{u}(z)},$
$\psi_{2}(z)=\frac{{\rm i}\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}+G\right)h(z)}{\vartheta\left(-\mathfrak{u}(z)+\mathfrak{u}(\infty)-\frac{\tau+1}{2}\right)}{\rm e}^{-{\rm i}\pi K\mathfrak{u}(z)}.$
These are essentially the same formulæ as (5.35), (5.36) except for a minor modification of the normalization. Accordingly, we define $\widehat{M}(z;G,K,\infty)$ exactly as in (5.37). The boundary relations for this matrix $\widehat{M}$ are now
$\widehat{M}(z_{+})=\widehat{M}(z_{-})\begin{bmatrix}0&1\\ -1&0\end{bmatrix},\qquad z\in\Gamma_{a_{+}}^{a_{-}},$
$\widehat{M}(z_{+})=\widehat{M}(z_{-})\begin{bmatrix}-{\rm e}^{-{\rm i}\pi K}&0\\ 0&-{\rm e}^{{\rm i}\pi K}\end{bmatrix},\qquad z\in\Gamma_{b}^{a_{+}},$
$\widehat{M}(z_{+})=\widehat{M}(z_{-})\begin{bmatrix}0&{\rm e}^{-2{\rm i}\pi(G-\frac{K\tau}{2})}\\ -{\rm e}^{2{\rm i}\pi(G-\frac{K\tau}{2})}&0\end{bmatrix},\qquad z\in\Gamma_{b}^{1}.$
By matching these to the boundary relations (5.44), we deduce the following.
Proposition 5.25.
The solution of the Riemann–Hilbert Problem 5.23 is given by $\widehat{M}(z;G,K)$ in (5.37) with the values of the constants $G$, $K$ given by
$G=\frac{1}{2{\rm i}\pi}\left(n\Omega_{1}+\nu+n\Omega_{2}\tau\right)+\frac{\tau}{2},\qquad K=\frac{n\Omega_{2}}{{\rm i}\pi}+1.$ (5.45)
The proof is entirely similar to Proposition 5.18. The same observation holds that the solvability of the Riemann–Hilbert Problem 5.23 and hence of the Riemann–Hilbert Problem 5.23 depends entirely on the non-vanishing of the expression
$\vartheta\left(\frac{n\Omega_{1}+\nu-n\Omega_{2}\tau}{2{\rm i}\pi}+\frac{1}{2}\right)$ (5.46)
@@
$\widetilde{H}_{1}:=\frac{n\Omega_{1}}{2{\rm i}\pi}+\operatorname{Re}\left(\frac{\nu}{2{\rm i}\pi}\right)-\frac{\operatorname{Re}\tau}{\operatorname{Im}\tau}\operatorname{Im}\left(\frac{\nu}{2{\rm i}\pi}\right)\in{\mathbb{Z}},$
$\widetilde{H}_{2}:=\frac{n\Omega_{2}}{2{\rm i}\pi}+\frac{\operatorname{Im}\left(\frac{\nu}{2{\rm i}\pi}\right)}{\operatorname{Im}\tau}+\frac{1}{2}\in{\mathbb{Z}}.$ (5.47)
The mesh of level sets $\widetilde{H}_{1}\in{\mathbb{Z}}\ni\widetilde{H}_{2}$ are the gridlines shown in Figure 3 in the portion $\operatorname{Re}(s)<0$ of each pane.
@@ 5.2.3 Local parametrices
The construction of the local parametrices is similar to that of Section 5.1.1. The only parametrix that is less commonly encountered (but see [10]) is the one that needs to be defined at the point $b$ in the case $\operatorname{Re}(s)>0$, when three main arcs meet at the same point (see Figure 8, left pane). Let $\mathbb{D}_{x}$ denote a small disk centered at $x\in\{a_{+},a_{-},b,1\}$ and not containing any of the other branch-points $_{+}, aa_{-}$, $b$, 11. Inside each of these disks one has to define an appropriate local solution of the Riemann–Hilbert Problem 5.12 or the Riemann–Hilbert Problem 5.20 (according to the sign of $\operatorname{Re}(s)>0$ or $\operatorname{Re}(s)<0$, respectively). Such a local solution (a.k.a. “local parametrix”) is constructed in terms of Airy functions or Bessel functions. More specifically,
  1. (1)
@@
  2. (2)
    Inside $\mathbb{D}_{b}$ the local parametrix is constructed also in terms of Airy functions, arranged however in a different formula along the lines explained in Appendix A.1.
  3. (3)
    Inside $\mathbb{D}_{1}$ the local parametrix is constructed in terms of Bessel functions. This can be found in (see [59, Section 3, p. 155–156]), or (see [8, Section 4.2]), and we will not enter into details since it is mainly irrelevant to the formulas that are of interest in this paper.
5.2.4 Summary: The approximating mesh of location of zeros
@@
$J_{\mathscr{E}}(z)=M_{{}_{Q}}(z)\left(\mathbf{1}+\mathcal{O}\bigl(n^{-1}\bigr)\right)M_{{}_{Q}}^{-1}(z).$
While in the region outside $\rm EoT$} the model solution $M_{{}_{Q}}$ does not depend on $n$ (see (5.7)) and hence $J_{\mathscr{E}}=\mathbf{1}+\mathcal{O}\bigl(n^{-1}\bigr)$ since the matrix $M_{{}_{Q}}$ is $n$-independent. On the other hand, when $s\in{\rm EoT}$}, the solution $M$ is instead given by Proposition 5.18 and consequently $M_{{}_{Q}}$ (5.26)) depends on $n$.
The validity of the estimate $J_{\mathscr{E}}=\mathbf{1}+\mathcal{O}\bigl(n^{-1}\bigr)$ depends on the matrices $M$, $M^{-1}$ remaining bounded, uniformly in $n$, with respect to $z$ on the contour. Given that $\det(M)\equiv 1$, the source of the potential unboundedness are the terms in the denominators in (5.37). Of these, the only term that may potentially vanish is the term $\vartheta\left(G-\frac{\tau+1}{2}\right)$, which becomes (5.40) or (5.46) (depending on the sign of $\operatorname{Re}(s)$). Therefore, the correct error estimate is
$J_{\mathscr{E}}(z)=\mathbf{1}+\mathcal{O}\left(\frac{1}{n\vartheta\left(G-\frac{\tau+1}{2}\right)^{2}}\right),$
with $G$ given by either (5.39) or (5.45) (depending on the sign of $\operatorname{Re}(s)$).
Thus the sufficient condition of asymptotic solvability of the Riemann–Hilbert Problem 5.2 is that
$\vartheta\left(\frac{n\Omega_{1}(s)-n\Omega_{2}(s)\tau(s)+\nu(s)}{2{\rm i}\pi}+\frac{1}{2}\mathcal{X}_{\{\operatorname{Re}s>0\}}(s)\right)=o\left(\frac{1}{\sqrt{n}}\right),$
with $\mathcal{X}_{\{\operatorname{Re}s>0\}}(s)$ denoting the indicator function of the right half-plane. This condition fails inside the union of small disks of radii $\mathcal{O}\bigl(n^{-\frac{3}{2}}\bigr)$ in the $s$-plane centered at the points where $\vartheta=0$}, namely, the intersection of the mesh of lines indicated by (5.41) (for $\operatorname{Re}(s)>0$) or (5.47) (for $\operatorname{Re}(s)<0$}).
The actual numerics shows a seemingly much faster convergence of the mesh to the actual position of the zeros (see Figure 3), which appears to be one of those serendipitous situations where the approximation works “better than expected”.
Figure 10: The RHPs for the error terms in the two cases $\operatorname{Re}(s)>0$ (left) and $\operatorname{Re}(s)<0$. Compare with Figure 8.
@@ 6 Approximation of the Hamiltonian
The Hamiltonian is obtained from the Taylor expansion near $z=0$ of the solution of Riemann–Hilbert Problem 5.2 outside of $\rm EoT$}, and Riemann–Hilbert Problem 5.10 inside $\rm EoT$}, according to Proposition 2.8.
In either case, the chain of transformations ($Y\to W\to T_{0}\sim M_{{}_{Q}}$) and the subsequent approximation imply that in a neighbourhood of $z=0$ we have
$Y_{n}(z)={\rm e}^{-n\frac{\ell}{2}\sigma_{3}}W(z){\rm e}^{n(g(z)+\frac{\ell}{2})\sigma_{3}}={\rm e}^{-n\frac{\ell}{2}\sigma_{3}}\mathscr{E}(z)M_{{}_{Q}}(z){\rm e}^{n(g(z)+\frac{\ell}{2})\sigma_{3}},$
where $\mathscr{E}(z)$, as discussed in Section 5.2.4, is a matrix that is $\mathbf{1}+\mathcal{O}\bigl(\frac{1}{n\vartheta(G-\frac{\tau+1}{2})^{2}}\bigr)$ near the origin (as long as $s$ has finite distance from the boundary of $\rm EoT$} and the imaginary axis within it).
@@
$Y_{n}^{-1}Y_{n}^{\prime}|_{z=0}=\bigl({\rm e}^{-ng\sigma_{3}}\bigl(M_{{}_{Q}}^{-1}M_{{}_{Q}}^{\prime}+M^{-1}_{{}_{Q}}\mathscr{E}^{-1}\mathscr{E}^{\prime}M_{{}_{Q}}\bigr){\rm e}^{ng\sigma_{3}}+ng^{\prime}(z)\sigma_{3}\bigr)\big|_{z=0}.$
Given the expression of $M$ in either Proposition 5.7 (for $s\not\in{\rm EoT}$), Proposition 5.18 for $s\in{\rm EoT}\cap\{\operatorname{Re}(s)>0\}$, or Proposition 5.25 $s\in{\rm EoT}\cap\{\operatorname{Re}(s)<0\}$, we have
$Y_{n}^{-1}Y_{n}^{\prime}|_{z=0}=\bigl({\rm e}^{(-ng-S)\sigma_{3}}\bigl(M^{-1}M^{\prime}+M^{-1}\mathscr{E}^{-1}\mathscr{E}^{\prime}M\bigr){\rm e}^{(S+ng)\sigma_{3}}+\bigl(S^{\prime}+ng^{\prime}(z)\bigr)\sigma_{3}\bigr)|_{z=0},$
@@
We are interested in the $(1,1)$ entry and hence the conjugation in the first term by the diagonal matrix ${\rm e}^{(S-ng)\sigma_{3}}$ is immaterial. Furthermore, the term containing $\mathscr{E}$ yields a sub-leading contribution which we ignore for the purpose of this computation. In principle, we should fork the computation according to the three regions: (i) outside $\rm EoT$}, (ii) inside $\rm EoT$} with $\operatorname{Re}(s)>0$, (iii) inside $\rm EoT$} with $\operatorname{Re}(s)<0$. However, the computation inside $\rm EoT$} would be rather formal because we have not specified the correct orders of approximations when $s$ is in a neighbourhood of one of the zeros. Additionally, the explicit approximation is of no particular interest to us and thus for simplicity we decided to forego it entirely in this paper.
Hamiltonian outside $\boldsymbol{{\rm EoT}}$. Using the explicit expression of $M(z)$ in Proposition 5.7, we have $M^{-1}M^{\prime}=\frac{{\rm i}\sigma_{2}}{s}$ and hence this term does not contribute being off-diagonal. Using (5.5) and the fact that $R(0)=-\frac{s}{2}$, $R(1)=-\sqrt{1+\smash{\frac{s^{2}}{4}}\vphantom{\big|}}$ with the root’s determination such that $R(1;s)\simeq-\frac{s}{2}$ as $|s|\to\infty$, we find $S^{\prime}(0)=\frac{{K}}{s}+\frac{2{\rho}}{s^{2}+s\sqrt{s^{2}+4}}$. Using then the explicit expression of $g(z)$ derived from (4.3)
$g^{\prime}(z)=\frac{1}{2}\varphi^{\prime}(z;s)+\frac{s}{2z^{2}}=\frac{\sqrt{z^{2}+\frac{s^{2}}{4}}}{z^{2}}+\frac{s}{2z^{2}}.$
Since $g^{\prime}(z)$ is regular at $z=0$, we need to take the root’s determination near $z=0$ that tends to $\frac{s}{2}$ and hence
$g^{\prime}(z)=-s\frac{\sqrt{1+\frac{4z^{2}}{s^{2}}}}{2z^{2}}+\frac{s}{2z^{2}}=-\frac{1}{s}+\mathcal{O}(z)\Rightarrow g^{\prime}(0;s)=-\frac{1}{s}.$
We thus conclude the following.
Proposition 6.1.
For $s$ in closed subsets outside $\rm EoT$}, we have the uniform approximation
$H_{V}(s)=\frac{n-{K}}{s}-\frac{2{\rho}}{s^{2}+s\sqrt{s^{2}+4}}+\frac{{\rho}}{2}+\mathcal{O}\bigl(n^{-1}\bigr).$